# polyBART: A Chemical Linguist for Polymer Property Prediction and Generative Design

### v1

@@ Abstract
Anagha Savit1, Harikrishna Sahu1, Shivank Shukla1, Wei Xiong1, Rampi Ramprasad1, 1 School of Materials Science and Engineering, Georgia Institute of Technology, Atlanta, Georgia, USA Correspondence: rampi.ramprasad@mse.gatech.edu
@@ 1 Introduction
Language models, owing to their generative capabilities, have emerged as an effective strategy for addressing both forward and inverse problems in the molecular space. Drawing inspiration from foundation models in Natural Language Processing (NLP), molecular foundation models (Edwards et al., 2022; Bagal et al., 2022) are pretrained on the ’language’ of molecules, typically represented as SMILES (Weininger, 1988), SMARTS (Daylight Chemical Information Systems, ), or SELFIES (Krenn et al., 2020) strings. These models are subsequently finetuned on downstream tasks, such as property prediction, hence solving the forward problem. The problem of generative design is often achieved by navigating the latent space learned during pretraining. The achievements of language models for molecules highlight a key opportunity for extending similar methodologies to polymers. Particularly, the structural and representational parallels between the molecular and polymer domains present a unique possibility of adapting existing molecular foundation models to the polymer space.
@@ 2.2 Polymer Informatics via NLP
ML algorithms have been applied in the polymer space for quite some time (Chen et al., 2021). Similar to molecules, a common approach in polymer informatics is to represent polymers numerically and use ML to connect these representations to points in the property space. A range of approaches, including regression algorithms (Doan Tran et al., 2020; Tao et al., 2021) and neural networks (Kuenneth et al., 2021b; Kuenneth et al., 2022) have been explored for this purpose. While the application of NLP in this space is relatively nascent, it has already seen significant success. In property prediction tasks, language models are typically used to generate numerical embeddings of polymer representations, which are then mapped to properties using approaches similar to traditional polymer informatics pipelines (Agarwal et al., 2025). Kuenneth and Ramprasad (2023) developed polyBERT by pretraining deBERTA (He et al., 2021) on millions of SMILES strings. Similar polymer language models have also been developed using alternative architectures, such as RoBERTa in TransPolymer (Xu et al., 2023) and T5 (Raffel et al., 2023) in PolyNC (Qiu et al., 2024).
For the inverse problem of generating polymer structures, early efforts have leveraged Variational Autoencoders (VAEs) (Batra et al., 2020). However, the lack of explicit chemical syntax in VAE architectures leads to invalid generations. By leveraging SELFIES (or PSELFIES in the present case), which guarantee validity, and language models that can capture structured sequences and syntactic patterns, we achieve state-of-the-art results in the generation of novel, valid, and property-aligned polymers.
@@ 3.1 PSELFIES
Table 1: Average RMSE and standard deviation over 5-fold CV test sets for the property prediction task. The best results are highlighted in bold and second-best are underlined.
@@ 4.1 Property Prediction
Baselines: We compare the performance of polyBART to the baseline SELFIES-TED model. Due to the absence of the [At] token in SELFIES-TEDsmall, our comparison is restricted to SELFIES-TEDlarge. In addition, we benchmark polyBART with polyBERT, current state-of-the-art transformer-based model for polymer property prediction. It is important to note that, given the novel bidirectional capabilities of polyBART, we are limited in comparison to unidirectional models developed for property prediction. To ensure a fair comparison of embedding quality, we use GPR across all models.
Results: Table 1 presents the performance of polyBARTsmall and polyBARTlarge, reporting the average RMSE across five CV test sets along with the corresponding standard deviations. We observe that our bidirectional polyBART performs on par with, and in many cases surpasses, polyBERT. Additionally, polyBARTsmall and polyBARTlarge outperform the baseline SELFIES-TEDlarge, achieving average RMSE reductions of 19.58% and 19.71%, respectively. These consistent performance gains underscore the ability of PSELFIES to specialize molecular language models to the polymer domain. Overall, our results confirm the capability of polyBART as a powerful polymer property predictor.
@@ 4.2 Impact of Transfer Learning
To further demonstrate the importance of PSELFIES in adapting molecular foundation models to the polymer space, we perform a comparative study. Specifically, we develop a model, BARTsmall, which shares the same architecture, training data, and configuration as polyBARTsmall, but is trained from scratch rather than initialized from SELFIES-TED. This setup allows us to isont the effect of transfer learning using PSELFIES.
@@ 4.3 Hypothetical Polymer Generation
polyBART’s generative capability, enabled by its decoder, is a key innovation of our work, setting it apart from existing polymer language models (Agarwal et al., 2025), which have been limited to unidirectional property prediction. To generate novel polymer structures, the PSELFIES sequence is first passed through the encoder to obtain a latent embedding. To explore the latent space, we apply Gaussian noise to this embedding using an n-fold sampling strategy: noise is added n times to generate n noised embeddings. These noisy embeddings are then decoded one by one using the polyBART decoder, and if a valid polymer is generated within these samples, it is retained. A grid search is used to tune the noise level for novelty and validity, details of which can be found in Appendix E.
We evaluate the generative capabilities of polyBART using a test set of 10,000 PSELFIES, all of which are unseen during the pretraining phase. We begin by examining the validity of the generated polymer structures. Specifically, we introduce two measures of validity: (i) molecule validity and (ii) polymer validity. Molecule validity checks whether the generated PSELFIES correspond to a chemically valid molecule, and is determined using RDKit. Polymer validity assesses whether the generated PSELFIES represent a structurally valid polymer by examining the placement and single-bond character of the [At] tokens. Since the [At] tokens denote the endpoints of the polymer repeat unit, a valid polymer must contain exactly two such tokens, each with a valency of one. The polymer validity metric enforces this requirement. Next, we assess the novelty of the valid generated structures, where a structure is considered novel if it passes both validity tests and is different from all input examples. Finally, we assess the quality of the generated structures using Fréchet ChemNet Distance (FCD) (Preuer et al., 2018) and Internal Diversity (IntDivp) metrics (Benhenda, 2017). FCD quantifies the similarity between the distributions of the generated and input structures based on the Euclidian distance between their respective embeddings. A higher FCD score indicates a greater divergence between the two distributions, suggesting that the generated structures differ from the input set. IntDivp evaluates the internal diversity of the generated structures, capturing the model’s tendency to produce structurally distinct outputs. We report both IntDiv1 (p = 1) and IntDiv2 (p = 2) scores, where higher values of these scores reflect greater diversity in the generated structures. Although these metrics were originally designed for molecules, by representing polymers as pseudo-molecules using PSELFIES, we can extend them to our case. See Appendix F for a detailed description of the metrics.
@@ 4.4 Property-conditioned Polymer Generation
Beyond general structure generation, a key capability of polyBART is its ability to generate novel polymers conditioned on desired target properties. At a high level, our approach enables property-guided generation by mapping the latent space of polymer embeddings to the property space and exploring regions with desired property values. We begin by identifying polymers that fall within a specified target range for a given property. These polymers are then passed through polyBART’s encoder to obtain embeddings. As in the general generation setting, we apply an n-fold Gaussian noise sampling strategy: each embedding is perturbed n times to generate n candidate embeddings, which are subsequently decoded and checked for validity. In this way, we are able to effectively explore neighborhoods in the latent space with desirable property values. The resulting candidate polymers are subjected to two filtering steps: (i) property filtering and (ii) Synthetic Accessibility (SA) score (Ertl and Schuffenhauer, 2009) filtering. To filter based on property, the generated candidates are passed through the GPR models developed for property prediction, and only those within a threshold range of the target property are retained. To assess synthesizability, we compute the SA score, a measure ranging from 1 (easy to make) to 10 (very difficult to make). For this, PSELFIES strings are converted to MSMILES, with the At atom replaced by hydrogen (H). The SA score is then calculated, and candidates with a score of 6 or lower are retained as viable candidates. Additional details are provided in Appendix G.
@@
As shown in Table 3, polyBARTsmall and polyBARTlarge are able to successfully generate hundreds of novel polymers that not only meet the desired property criteria but are also predicted to be synthetically accessible. Examples of novel structures generated by our models are illustrated in Figure 3. Additionally, visualizations of the property-conditioned latent space can be found in Appendix H.
@@ 6 Conclusion
In this work, we propose PSELFIES, a novel representation of polymers that enables the transfer of learned knowledge from molecular language models to the polymer space. We develop polyBARTsmall and polyBARTlarge, two polymer foundation models created through the continued pretraining of SELFIES-TEDsmall and SELFIES-TEDlarge, respectively. Our results show that polyBART achieves excellent performance in property prediction. Moreover, polyBART uniquely combines property prediction with generative design in a single unified framework and consistently produces numerous novel polymer structures. Generated candidates not only satisfy target property criteria but also exhibit synthetic accessibility. By successfully synthesizing one such predicted polymer and experimentally confirming our predictions, we illustrate that language models can reliably drive real-world advances in polymer design. Our work enables true bidirectional translation between polymer structures and properties and highlights the potential of language models in advancing polymer informatics.
@@ 8 Acknowledgements
The authors thank the Office of Naval Research (ONR) for financial sponsorship of this work through Grant N00014-23-1-2279. The authors are also grateful for the usage of the Gutekunst lab for synthesis and characterization of one of the designed polymers,
@@ Appendix D Property Datasets
We evaluate polyBART on Tg, Td, Tm, Egc, Egb, Eea, and Ei. Histograms of the value distributions for each property are shown below. Thermal properties (Tg, Td, Tm) are obtained from PolyInfo at https://polymer.nims.go.jp/. Egc and Egb are taken from the polyVERSE repository (https://github.com/Ramprasad-Group/polyVERSE/tree/main), Eea and Ei are obtained from the Khazana polymer database (https://khazana.gatech.edu/dataset/).
@@ D.2 Electronic Properties
Appendix E Optimal Noise Level
@@ F.2 IntDivp
Appendix G Threshold Selection for SA Scores
@@
Appendix H Latent Space Visualization
@@
Appendix I Experimental Procedure

### v2

@@ Abstract
Anagha Savit1, Harikrishna Sahu1, Shivank Shukla1, Wei Xiong1, Rampi Ramprasad1 1 School of Materials Science and Engineering, Georgia Institute of Technology Correspondence: rampi.ramprasad@mse.gatech.edu
@@ 1 Introduction
Language models, owing to their generative capabilities, have emerged as an effective strategy for addressing both forward and inverse problems in the molecular space. Drawing inspiration from foundation models in Natural Language Processing (NLP), molecular foundation models (Edwards et al., 2022; Bagal et al., 2022) are pretrained on the ’language’ of molecules, typically represented as SMILES (Weininger, 1988), SMARTS (Daylight Chemical Information Systems, ), or SELFIES (Krenn et al., 2020) strings. These models are subsequently fine-tuned on downstream tasks, such as property prediction, hence solving the forward problem. The problem of generative design is often achieved by navigating the latent space learned during pretraining. The achievements of language models for molecules highlight a key opportunity for extending similar methodologies to polymers. Particularly, the structural and representational parallels between the molecular and polymer domains present a unique possibility of adapting existing molecular foundation models to the polymer space.
@@ 2.2 Polymer Informatics via NLP
ML algorithms have been applied in the polymer space for quite some time (Chen et al., 2021). Similar to molecules, a common approach in polymer informatics is to represent polymers numerically and use ML to connect these representations to points in the property space. A range of approaches, including regression algorithms (Doan Tran et al., 2020; Tao et al., 2021) and neural networks (Kuenneth et al., 2021b; Kuenneth et al., 2022) have been explored for this purpose. Graph neural networks (GNNs) are also widely used in polymer modeling due to their ability to capture both local and global topological features of polymer graphs. Gurnani et al. (2023) developed polyGNN, a multitask polymer GNN model for polymer property prediction, demonstrating strong performance across numerous properties. While the application of NLP in this space is relatively nascent, it has already seen significant success. In property prediction tasks, language models are typically used to generate numerical embeddings of polymer representations, which are then mapped to properties using approaches similar to traditional polymer informatics pipelines (Agarwal et al., 2025). Kuenneth and Ramprasad (2023) developed polyBERT by pretraining deBERTA (He et al., 2021) on millions of SMILES strings. Similar polymer language models have also been developed using alternative architectures, such as RoBERTa in TransPolymer (Xu et al., 2023) and T5 (Raffel et al., 2023) in PolyNC (Qiu et al., 2024).
For the inverse problem of generating polymer structures, early efforts have leveraged Variational Autoencoders (VAEs) (Batra et al., 2020). However, the lack of explicit chemical syntax in VAE architectures leads to invalid generations. Transformers can learn context-aware representations of molecular structures (Dollar et al., 2021), with each attention head learning to focus on distinct substructures. By leveraging SELFIES (or PSELFIES in the present case), which guarantee validity, and language models that can capture structured sequences and syntactic patterns, we achieve state-of-the-art results in the generation of novel, valid, and property-aligned polymers.
@@ 3.1 PSELFIES
LLaMA-3-8B 58.13 95.28 71.91 - - - -
Table 1: Performance on the property prediction tasks. The best results are highlighted in bold and second-best are underlined.
@@ 4.1 Property Prediction
Baselines: We compare the performance of polyBART to the baseline SELFIES-TED model. Due to the absence of the [At] token in SELFIES-TEDsmall, our comparison is restricted to SELFIES-TEDlarge. In addition, we benchmark polyBART with polyBERT, current state-of-the-art transformer-based model for polymer property prediction. It is important to note that, given the novel bidirectional capabilities of polyBART, we are limited in comparison to unidirectional models developed for property prediction. To ensure a fair comparison of embedding quality, we use GPR across all models. We additionally compare polyBART with LLaMA, with details regarding the fine-tuning experiment given in Appendix E.
Results: Table 1 presents the performance of polyBARTsmall and polyBARTlarge, reporting the average RMSE across five CV test sets along with the corresponding standard deviations. We observe that our bidirectional polyBART performs on par with, and in many cases surpasses, polyBERT. Notably, while matching the predictive accuracy of the unidirectional polyBERT, the bidirectional polyBART offers the added advantage of generating entirely new polymer candidates not present in the training data. Additionally, polyBARTsmall and polyBARTlarge outperform the baseline SELFIES-TEDlarge, achieving average RMSE reductions of 19.4% and 19.7%, respectively. These consistent performance gains underscore the ability of PSELFIES to specialize molecular language models to the polymer domain. We also notice that polyBART outperforms LLaMA-3 on thermal property prediction. Overall, our results confirm the capability of polyBART as a powerful polymer property predictor.
@@ 4.2 Impact of Transfer Learning
To further demonstrate the importance of PSELFIES in adapting molecular foundation models to the polymer space, we perform a comparative study. Specifically, we develop a model, BARTsmall, which shares the same architecture, training data, and configuration as polyBARTsmall, but is trained from scratch rather than initialized from SELFIES-TED. This setup allows us to isolate the effect of transfer learning using PSELFIES.
@@ 4.3 Hypothetical Polymer Generation
polyBART’s generative capability, enabled by its decoder, is a key innovation of our work, setting it apart from existing polymer language models (Agarwal et al., 2025), which have been limited to unidirectional property prediction. To generate novel polymer structures, the PSELFIES sequence is first passed through the encoder to obtain a latent embedding. To explore the latent space, we apply Gaussian noise to this embedding using an n-fold sampling strategy: noise is added n times to generate n noised embeddings. These noisy embeddings are then decoded one by one using the polyBART decoder, and if a valid polymer is generated within these samples, it is retained. A grid search is used to tune the noise level for novelty and validity, details of which can be found in Appendix F.
We evaluate the generative capabilities of polyBART using a test set of 10,000 PSELFIES, all of which are unseen during the pretraining phase. We begin by examining the validity of the generated polymer structures. Specifically, we introduce two measures of validity: (i) molecule validity and (ii) polymer validity. Molecule validity checks whether the generated PSELFIES correspond to a chemically valid molecule, and is determined using RDKit. Polymer validity assesses whether the generated PSELFIES represent a structurally valid polymer by examining the placement and single-bond character of the [At] tokens. Since the [At] tokens denote the endpoints of the polymer repeat unit, a valid polymer must contain exactly two such tokens, each with a valency of one. The polymer validity metric enforces this requirement. Next, we assess the novelty of the valid generated structures, where a structure is considered novel if it passes both validity tests and is different from all input examples. Finally, we assess the quality of the generated structures using Fréchet ChemNet Distance (FCD) (Preuer et al., 2018) and Internal Diversity (IntDivp) metrics (Benhenda, 2017). FCD quantifies the similarity between the distributions of the generated and input structures based on the Euclidian distance between their respective embeddings. A higher FCD score indicates a greater divergence between the two distributions, suggesting that the generated structures differ from the input set. IntDivp evaluates the internal diversity of the generated structures, capturing the model’s tendency to produce structurally distinct outputs. We report both IntDiv1 (p = 1) and IntDiv2 (p = 2) scores, where higher values of these scores reflect greater diversity in the generated structures. Although these metrics were originally designed for molecules, by representing polymers as pseudo-molecules using PSELFIES, we can extend them to our case. See Appendix G for a detailed description of the metrics.
@@ 4.4 Property-conditioned Polymer Generation
Beyond general structure generation, a key capability of polyBART is its ability to generate novel polymers conditioned on desired target properties. At a high level, our approach enables property-guided generation by mapping the latent space of polymer embeddings to the property space and exploring regions with desired property values. We begin by identifying polymers that fall within a specified target range for a given property. These polymers are then passed through polyBART’s encoder to obtain embeddings. As in the general generation setting, we apply an n-fold Gaussian noise sampling strategy: each embedding is perturbed n times to generate n candidate embeddings, which are subsequently decoded and checked for validity. In this way, we are able to effectively explore neighborhoods in the latent space with desirable property values. The resulting candidate polymers are subjected to two filtering steps: (i) property filtering and (ii) Synthetic Accessibility (SA) score (Ertl and Schuffenhauer, 2009) filtering. To filter based on property, the generated candidates are passed through the GPR models developed for property prediction, and only those within a threshold range of the target property are retained. To assess synthesizability, we compute the SA score, a measure ranging from 1 (easy to make) to 10 (very difficult to make). For this, PSELFIES strings are converted to MSMILES, with the At atom replaced by hydrogen (H). The SA score is then calculated, and candidates with a score of 6 or lower are retained as viable candidates. Additional details are provided in Appendix H.
@@
As shown in Table 3, polyBARTsmall and polyBARTlarge are able to successfully generate hundreds of novel polymers that not only meet the desired property criteria but are also predicted to be synthetically accessible. Examples of novel structures generated by our models are illustrated in Figure 3. Additionally, visualizations of the property-conditioned latent space can be found in Appendix I.
@@ 6 Conclusion
In this work, we propose PSELFIES, a novel representation of polymers that enables the transfer of learned knowledge from molecular language models to the polymer space. We develop polyBARTsmall and polyBARTlarge, two polymer foundation models created through the continued pretraining of SELFIES-TEDsmall and SELFIES-TEDlarge, respectively. Our results show that polyBART performs comparably to state-of-the-art approaches in property prediction. Moreover, polyBART uniquely combines property prediction with generative design in a single unified framework and consistently produces numerous novel polymer structures. Generated candidates not only satisfy target property criteria but also exhibit synthetic accessibility. By successfully synthesizing one such predicted polymer and experimentally confirming our predictions, we illustrate that language models can reliably drive real-world advances in polymer design. Our work enables true bidirectional translation between polymer structures and properties and highlights the potential of language models in advancing polymer informatics.
@@ 8 Acknowledgements
The authors thank the Office of Naval Research (ONR) for financial sponsorship of this work through Grant N00014-23-1-2279. The authors are also grateful for the usage of the Gutekunst lab for synthesis and characterization of one of the designed polymers.
@@ Appendix D Property Datasets
We evaluate polyBART on Tg, Td, Tm, Egc, Egb, Eea, and Ei. Histograms of the value distributions for each property are shown below. Thermal properties are obtained from PolyInfo at https://polymer.nims.go.jp/. Electronic properties are accessible through the polyVERSE repository (https://github.com/Ramprasad-Group/polyVERSE/tree/main) and the Khazana polymer database (https://khazana.gatech.edu/dataset/).
@@ D.2 Electronic Properties
Appendix E Comparison with LLaMA
To enable a comparison between polyBART and an LLM, we evaluate both approaches on the task of thermal property prediction, using the 8B-parameter LLaMA-3 model as the representative LLM baseline. Because the available electronic-property datasets are comparatively small, we limit this comparison to thermal properties.
We fine-tune LLaMA to follow a simple question–answer protocol that maps SELFIES to scalar property values. Specifically, each training instance is formatted as:
Prompt: What is the {property name} of the polymer with SELFIES: {selfies}?
Response: {property value}
We fine-tune LLaMA with LoRA ($r=8$, $\alpha=8$) and early stopping. At inference time, we employ deterministic decoding (temperature =0=0, no sampling) to produce a single numeric answer.
Appendix F Optimal Noise Level
@@ G.2 IntDivp
Appendix H Threshold Selection for SA Scores
@@
Appendix I Latent Space Visualization
@@
Appendix J Experimental Procedure